# thanks: Contributed equallythanks: Contributed equally

### v3

@@ Abstract
Collective motion as a flock is a widely observed phenomenon in active matter systems. We report a flocking transition mechanism in a system of chemically interacting active colloidal particles sustained purely by chemo-repulsive torques at low to medium densities. The basic requirements to maintain the polar liquid flock are excluded volume repulsions and deterministic long-ranged net repulsive torques. The mechanism we report requires that the time scale individual colloids move a unit length to be dominant with respect to the time they deterministically sense chemicals. This can be equivalently interpreted as pair colloids sliding a minimal unit length before deterministically rotating due to chemical interactions. Switching on the translational repulsive forces renders the flock a crystalline structure. We complement these results with an analysis of a continuum hydrodynamic model, with the transition corresponding to destabilization of the flocking state.
@@ I Introduction
The collective swarming behaviour of interacting agents - also called flocking [1] - has been a topic of sustained interest in the field of non-equilibrium statistical physics. Flocking is also a ubiquitous phenomena in the natural world, widely reported in groups of animals such as birds, fish, and mammals [2]. Paradigmatic models that incorporate local alignment interactions to study flocking have been extensively studied [3, 4, 5, 1] and continue to be investigated [6, 7, 8, 9]. In the Vicsek model [3], local alignment interactions among the individual agents can lead to transition from a disordered state to one with large-scale, coordinated movement. The flocking transition in Vicsek-like models with short-range interactions has been widely studied, and traveling density bands are observed at the onset of transition [4, 5, 10, 11].
@@
The possibility of attaining an emergent global polar order (flocking) arising purely out of dynamical particle-based models, without any explicit alignment interaction, is a subject of recent interest [12]. Studies on these include those studying the interplay between non-equilibrium phase separation [13] and other collective dynamics [14, 15, 16, 17], those that study the effect of velocity alignment interactions [18, 19, 20], models incorporating attractive interactions [21, 22], history-dependence [23, 24], and those that exclusively study the effect of repulsive forces and torques [25, 26]. Though various possible mechanisms deviating from the Vicsek-like phenomenology have been established, the explicit relevance of the separate force and torque contributions have not been established. For instance, migrating cells that display a global polarity have been known to display chemo-tactic attractive interactions in addition to (among others) either a contact “inhibited” or “attracted” locomotion - akin to a long-ranged torque of either repulsive or attractive nature [27, 28, 22]. Thus, open questions remain on the minimal ingredients and mechanism at play arising from the rotational and translational components separately.
@@
In this paper, we report a minimal mechanism for the formation of polar liquids where colliding pair particles have to sufficiently slowly rotate in response to chemical gradients before moving a unit distance upon collision. The basic ingredients that are sufficient to produce such a flock are thus (i) short-ranged excluded volume repulsion, and (ii) long-ranged (chemo) repulsive torques. There is no need for any additional (chemo)repulsive translational forces between the particles, nor is there a requisite role for noise in destabilizing the flock. Furthermore, the incorporation of long-ranged translational chemo-repulsive forces between the particles induces instead a crystalline flock (with regular lattice spacing). Our work thus adds insights to investigations on attaining a global polar order via dynamical mechanisms (i.e. without any explicit alignment interaction) by providing a minimal mechanism, both for flock formation and destabilization, that, to our best knowledge, has not been reported elsewhere.
The remainder of the paper is organized as follows. In section II, we describe our model to study phoretic (chemical) interactions of active particles. In section III, we present our results from particle-based numerical simulations by mapping phase diagrams and describing the properties of the two distinct flocking states. The two distinct flocking states are termed as: CLF (Chemorepulsive Liquid Flocks), and CCF (Chemorepulsive Crystalline Flocks). Chemorepulsive torques between the particles, which turn them away from each other [15, 23, 26], are essential for the formation of both CLF and CCF states. For the formation of CCF, we require both chemorepulsive force as well as chemorepulsive torques. See below for details. Then, we complement the results from particle-based simulation with a stability analysis of a continuum description in section IV. Finally, we conclude in section V by summarizing our main results and comparing it with existing literature, along with suggesting directions for future investigations.
Figure 1: Polar liquid flocks for $\chi_{t}=0$. (a) Phase diagram in the $(\rho,l_{r}/b)$ plane. Here, $l_{-}/b$ is annotated as a dotted red line, while the density dependent $l_{+}/b$ in dashed black line. Symbols here and throughout the paper denoting a flock (yellow triangle), bands (green circles), and disordered (blue squares). These are distinguished with different values of polar order parameter $\mathbf{P}$: flock ($\mathbf{P}>0.9$), bands ($0.1<\mathbf{P}<0.9$) and disordered ($\mathbf{P}<0.1$). Marked (I)-(III) scenario and transition lines (black and red) are described in the schematic shown in Fig. 2. Snapshots of the respective phases are shown in (b)-(d) respectively (color coding for orientation of each particle is in middle panel). (e) The evolution of the polarization is displayed for the corresponding states (similarly colored). See Movie I [29] for dynamics to obtain CLF starting with a disordered state.
Figure 2: Schematic of the mechanism for formation of fluid flocks. Top-to-bottom: Series of scenarios depicted corresponding to parameters choices labelled in Fig.1a. Top (I) and bottom (III) panels show scenarios where the chemical response is either too slow or too quick, such that a stable flock cannot be formed. The middle panel (II) shows the pair collision mechanism that would produce a flock if the time scale follow: $\tau_{sl}^{+}<\tau_{sl}<\tau_{sl}^{-}$. See text for details.
@@ II Model
We consider a set of $N$ chemically interacting polar colloids. We model the $i$th active particle as a colloid particle centered at $\mathbf{r}_{i}=(x_{i},y_{i})$, confined to move in two-dimensions, which self-propels with a speed $v_{0}$, along the directions $\mathbf{e}_{i}=(\cos\theta_{i},\,\sin\theta_{i})$. Here $i=1,2,3,\dots,N$ and $\theta_{i}$ is the angle made by the orientation $\mathbf{e}_{i}$ with respect to the positive $x$-axis. The orientation $\mathbf{e}_{i}$ of the $i$th particle, given in terms of the angle $\theta_{i}$, changes due to coupling its dynamics to the dynamics of a phoretic (scalar) field $c$, as we describe below. The position $\mathbf{r}_{i}$ and orientation $\mathbf{e}_{i}$ of the $i$th particle is updated as:
$\dot{\mathbf{r}}_{i}={\mathbf{V}}_{i},\qquad\dot{\mathbf{e}}_{i}$ $={\mathbf{\Omega}}_{i}\times\mathbf{e}_{i},$ (1)
Here, the translational velocity $\mathbf{V}_{i}$ and angular velocity $\mathbf{\Omega}_{i}$ of the $i$th particle are given as:
${\mathbf{V}}_{i}$ $=v_{0}\mathbf{e}_{i}+\chi_{t}\,\mathbf{J}_{i}+\mu\mathbf{F}_{i}+\bm{\eta}^{t}_{i},$ (2a)
$\qquad\mathbf{\Omega}_{i}$ $=\chi_{r}\left(\mathbf{e}_{i}\times\mathbf{\mathbf{{J}}}_{i}.\right)+\bm{\eta}^{r}_{i}.$ (2b)
@@
Here $v_{0}$ is the self-propulsion speed of an isolated active particle. The term $\mathbf{J}_{i}$ is chemical flux on the location of the $i$th particle, which gives the inter-particle interaction (of chemical origin). The phoretic flux is given as: $\mathbf{\mathbf{\mathbf{{J}}}}_{i}(t)=-\left[\mathbf{\nabla}c(\mathbf{r},t)\right]_{\mathbf{r}=\mathbf{r}_{i}}$, where $c(\mathbf{r},t)$ is the concentration of the chemicals (e.g filled micelles in an oil-emulsion system [30, 23]). The concentration field evolves in time as:
@@
Here, $D_{c}$ is the diffusion coefficient and $c_{0}$ is emission constant of the chemical. We assume the limit in which chemical field leads to instantaneous interactions between the colloidal surfaces [31, 32, 33, 34]. The solution of the Eq.3 can be used to compute $\mathbf{{J}}_{i}$ as:
$\mathbf{J}_{i}=\frac{c_{0}}{4\pi D_{c}}\sum^{N}_{\begin{subarray}{c}j=1\\ i\neq j\end{subarray}}\frac{\mathbf{r}_{ij}}{r_{ij}^{3}}$ (4)
Here, $r_{ij}=|\mathbf{r}_{ij}|$, with $\mathbf{r}_{ij}=\mathbf{r}_{i}-\mathbf{r}_{j}$. The term proportional to $\chi_{r}$ in Eq.(2) drives orientational changes (turning away from each other if $\chi_{r}>0$) through interparticle chemical interactions. In this paper, we only consider case of $\chi_{r}>0$ when particles turn away from each other due to chemical interactions [24, 23]. Similarly, the term proportional to $\chi_{t}$ ensures repulsion in the positional dynamics if $\chi_{t}>0$. Here, both $\chi_{r}$ and $\chi_{t}$ are positive (unless otherwise stated) and the system is said to be chemo-repulsive.
In Eqs.(1)-(2), to preclude overlap of the particles, there is a force on the particles: $\mathbf{F}_{i}=-\boldsymbol{\nabla}_{i}\mathcal{U}$, while, $\mathcal{U}=\sum_{i<j}\mathcal{U}^{e}(\mathbf{r}_{i},\mathbf{r}_{j}).$ With $b$ as the colloidal radius, we choose $\mathcal{U}^{e}$ to be of the form: $\mathcal{U}^{e}=\kappa\left(r_{ij}-2b\right)^{2}$ if $r_{ij}<2b$, while it vanishes otherwise. Here, $\kappa$ is a constant. In Eq.2, $\bm{\eta}^{t}_{i}$ and $\bm{\eta}^{r}_{i}$ are white noises with zero mean and no temporal correlation. The variance of noises $\bm{\eta}^{t}_{i}$ and $\bm{\eta}^{r}_{i}$, are respectively, $2D_{t}$ and $2D_{r}$. These noises terms are included for the sake of completion and to demonstrate that our results are robust against weak fluctuations. In this paper, the fluctuations are set to be negligibly small to emphasize that flocks can be formed and destroyed by deterministic causes  [^1] Indeed, the polar order is expected to be destroyed at larger values of fluctuations [3]. In this paper, we choose noise to be less dominant to deterministic effects by choosing Péclet number $\mathrm{Pe}=v_{0}/(bD_{r})$ to be around $10^{9}$. Simulation details along with the full set of parameters used to generate each figure is given in the Appendix A. The dimensions of chemical concentration is: $c]=[\text{L}^{-3}]$, and chemical flux are $\mathbf{J}]=[\text{L}^{-4}]$. Thus, the dimensions of parameters are $\chi_{r}]=[\text{L}^{4}\text{T}^{-1}]$ and $\chi_{t}]=[\text{L}^{5}\text{T}^{-1}]$.
III Results
We first summarize our main results for the case of $\chi_{t}=0$ (no chemorepulsive forces between the particles). In particular, we report here polar fluids that are obtained due to chemo-repulsive torques alone ($\chi_{r}>0$). See Fig.1. These flocks are obtained only if pair colloids move at least a distance ($l_{sl}^{+}$, defined later) before deterministically rotating away due to sensing of chemical gradients. The schematic of the mechanism for polar fluids that we report is presented in Fig. 2. For $\chi_{t}>0$, we obtain crystalline flocks that are obtained via a combination of repulsive torques and forces (Fig. 3(c)). We also find these crystalline flocks for (a narrow range of) attractive forces (Fig. 3(b)). In this paper, we denote the former as “Chemorepulsive Liquid Flocks” (CLF), and the later as “Chemorepulsive Crystalline Flocks” (CCF).
Figure 3: (a) Phase diagram in the $(l_{r}/b,l_{t}/b)$ plane for the case of $\chi_{t}>0$. Dashed line refers to lower bound on $\chi_{r}$, solid line for lower bound on $\chi_{r}$ from Eq. (8). (b) Similar phase diagram for the case of $\chi_{t}<0$ (extension of (a)). The transition line is extrapolated from (3)(a). (c) Shows a snapshot of the CCF phase (same color scheme as in Fig (1)). See Movie II [29] for dynamics to obtain CCF starting with a disordered state. (d) and (e) compare $g(r,\phi)$ for the case of the CLF and the CCF. In (f), a comparison of the flocking phase, for CLF and CCF, is shown for the pair correlation $g(r)$.
III.1 CLFs via long-ranged repulsive torques

### v4

@@ Abstract
Coherent collective motion is a widely observed phenomenon in active matter systems. Here, we report a flocking transition mechanism in a system of chemically interacting active colloidal particles sustained purely by chemo-repulsive torques at low to medium densities. The basic requirements to maintain the global polar order are excluded volume repulsions and long-ranged repulsive torques. This mechanism requires that the time scale individual colloids move a unit length to be dominant with respect to the time they deterministically respond to chemical gradients, or equivalently, pair colloids sliding together a minimal unit length before deterministically rotating away from each other. Switching on the translational repulsive forces renders the flock a crystalline structure. Furthermore, liquid flocks are observed for a range of chemo-attractive inter-particle forces. Various properties of these two distinct flocking phases are contrasted and discussed. We complement these results with stability analysis of a hydrodynamic model, which admits the transition corresponding to destabilization of the flocking state observed in particle-based simulations.
@@ I Introduction
The collective swarming behaviour of interacting agents - also called flocking [1] - has been a topic of sustained interest in the field of non-equilibrium statistical physics. Flocking is also a ubiquitous phenomena in the natural world, widely reported in groups of animals, such as birds, fish, and mammals [2, 3, 4]. Paradigmatic models that incorporate local alignment interactions to study flocking have been extensively studied [5, 6, 7, 8, 1] and continue to be investigated [9, 10, 11, 12]. Most notably, in the Vicsek model [5], local alignment interactions among the individual agents can lead to transition from a disordered state to one with large-scale, coordinated movement. The flocking transition in these models with short-range alignment interactions has been widely studied, along with extensive investigation on the nature of the order-to-disorder transition [6, 8, 13, 14] and universality of critical exponents [1].
The possibility of attaining an emergent global polar order (flocking) arising purely out of dynamical particle-based models, without any explicit alignment interaction, is a subject of recent interest, and recently reviewed in [15]. Systems where such a phase has been observed are typically over-damped [16, 17, 18] and include those with velocity alignment interactions [17, 19, 20, 21], distance-dependent interactions [22], history-dependence [23, 24], and those with repulsive torques and forces between the particles without volume exclusion [25, 26]. A separate class of inertial models have also reported such a transition [27, 28, 29, 30], including those with attractive interactions [31].
@@
Though various possible mechanisms deviating from the Vicsek-like rule have been established, the explicit relevance of the long-ranged interparticle forces and torques between active particles have not been established. For instance, migrating cells that display a global polarity have been known to display chemo-tactic attractive interactions in addition to (among others) either a contact “inhibited” or “attracted” locomotion - akin to a long-ranged torque of either repulsive or attractive nature [32, 33, 18]. Studies on translational and rotational diffusiophoretic motion of active particles (with self-generated chemical field), with attractive torques results collapse of particles into dense clusters [34]. However, extensive study on repulsive torque is rare in these systems. Thus, open questions remain on the minimal ingredients and mechanism at play arising from the forces and torques of inter-particle phoretic interactions.
@@
In this paper, we report a minimal mechanism for the formation of liquid flocks, which we herein denote as “Chemorepulsive Liquid Flocks” (CLF). The basic ingredients that are sufficient to produce such a flock are: (i) short-ranged excluded volume repulsion, and (ii) long-ranged (chemo) repulsive torques. There is no need for any additional (chemo)repulsive translational forces between the particles, nor is there a requisite role for noise in destabilizing the flock. Furthermore, the incorporation of long-ranged translational chemo-repulsive forces between the particles induces instead a crystalline flock (with regular lattice spacing), we denote this herein as “Chemorepulsive Crystalline Flocks” (CCF). Our work thus adds insights to investigations on attaining a global polar order via dynamical mechanisms without any explicit alignment interaction, by providing a minimal mechanism, both for flock formation and destabilization, that, to our best knowledge, has not been reported elsewhere.
The remainder of the paper is organized as follows. In Section II, we describe our model to study phoretic (chemical) interactions of active particles, and enumerate various important dimensionless parameters of our study. In Section III, we report various simulation results for the CLF, and likewise for the CCF in section IV. A comparison of these two phases is done in section V by studying the pair correlation, hexatic order, order-to-disorder transition, density fluctuations and finite-size effects. In Section VI, we demonstrate that flocking persists even for chemo-attractive forces between particles which turn away from each other. Further, we complement the results from particle-based simulation with a stability analysis of a continuum model of coupled density and polarization fields in Section VII. Finally, conclusions and a discussion of our work in relation to other findings is presented in Section VIII.
@@ II Model
The model consists of two dynamical parts: (a) the dynamics of particles, and (b) the dynamics of the chemical concentration field; these are described in the following subsections.
II.1 Dynamics of the particles
@@
We consider a set of $N$ chemically interacting active colloids of radius $b$. We model the $i$th active particle as a colloid particle centered at $\mathbf{r}_{i}=(x_{i},y_{i})$, confined to move in two-dimensions, which self-propels with a speed $v_{0}$, along the directions $\mathbf{e}_{i}=(\cos\theta_{i},\,\sin\theta_{i})$. Here $i=1,2,3,\dots,N$ and $\theta_{i}$ is the angle made by the orientation $\mathbf{e}_{i}$ with respect to the positive $x$-axis. The orientation $\mathbf{e}_{i}$ of the $i$th particle, given in terms of the angle $\theta_{i}$, changes due to coupling its dynamics to the dynamics of a phoretic (scalar) field $c$, as we describe below. The position $\mathbf{r}_{i}$ and orientation $\mathbf{e}_{i}$ of the $i$th particle is updated as:
$\dot{\mathbf{r}}_{i}$ $=v_{0}\mathbf{e}_{i}+\chi_{t}\,\mathbf{J}_{i}+\mu\mathbf{F}_{i}+\bm{\eta}^{t}_{i},$ (1a)
$\qquad\dot{\mathbf{e}}_{i}$ $=\chi_{r}\Big[\left(\mathbf{e}_{i}\times\mathbf{\mathbf{{J}}}_{i}\right)+\bm{\eta}^{r}_{i}\Big]\times\mathbf{e}_{i}.$ (1b)
@@
Here, $v_{0}$ is the self-propulsion speed of an isolated active particle. The terms $\bm{\eta}^{t}_{i}$ and $\bm{\eta}^{r}_{i}$ are white noises with zero mean and no temporal correlation. The variance of noises $\bm{\eta}^{t}_{i}$ and $\bm{\eta}^{r}_{i}$, are respectively, $2D_{t}$ and $2D_{r}$. These noises terms are included for the sake of completion and to demonstrate that the results are robust against weak fluctuations. The term $\mathbf{J}_{i}$ is chemical flux on the location of the $i$th particle, which gives the inter-particle interaction (of chemical origin). The phoretic flux is given as: $\mathbf{\mathbf{\mathbf{{J}}}}_{i}(t)=-\left[\mathbf{\nabla}c(\mathbf{r},t)\right]_{\mathbf{r}=\mathbf{r}_{i}}$, where $c(\mathbf{r},t)$ is the concentration of the chemicals (e.g filled micelles in an oil-emulsion system [35, 23, 36]).
The term proportional to $\chi_{r}$ in Eq.(1) drives orientational changes (turning particles away from each other if $\chi_{r}>0$) through interparticle chemical interactions. In this paper, we only consider case of $\chi_{r}>0$ when particles turn away from each other due to chemical interactions [24, 23]. Similarly, the term proportional to $\chi_{t}$ ensures repulsion in the positional dynamics if $\chi_{t}>0$. If both $\chi_{r}$ and $\chi_{t}$ are positive, then the system is said to be chemo-repulsive. A schematic representation of the dynamics due to the model is given in Fig.1 to described the physical meaning of parameters $\chi_{r}$ (which controls rotation) and $\chi_{t}$ (which controls translation) for chemical interactions between the particles.
Figure 1: Schematic description of the chemical interactions between the particles based on Eq.(1). The black dot on the orange circles indicate the orientation of the particles. Black curved arrows show rotations from chemical interactions, while solid red arrows shows translations from chemical interactions. The particles turn away from each other if $\chi_{r}>0$, while they turn towards each other if $\chi_{r}<0$. Similarly, particles translate away from each other if $\chi_{t}>0$, while they translate towards each other if $\chi_{t}<0$.
In Eq.(1), to preclude overlap of the particles, there is a short-ranged purely repulsive body force $\mathbf{F}_{i}$ on the particles. The expression of the body force on the $i$th particle is given as:
$\mathbf{F}_{i}=-\frac{\partial U}{\partial\mathbf{r}_{i}},\qquad U=\sum_{i<j}\mathcal{U}^{e}(\mathbf{r}_{i},\mathbf{r}_{j}).$ (2)
With $b$ as the colloidal radius and $r_{ij}=|\mathbf{r}_{i}-\mathbf{r}_{j}|$, we choose $\mathcal{U}^{e}$ to be of the form:
$\mathcal{U}^{e}=\begin{cases}\kappa\left(r_{ij}-2b\right)^{2},\qquad\quad\text{if }r_{ij}<2b,\\ 0,\qquad\qquad\qquad\qquad\text{otherwise.}\end{cases}$
Here, $\kappa$ is a constant, which determines the strength of the excluded volume repulsion.
II.2 Dynamics of the chemical field
Though the focus of this paper is on chemical interactions, we note that our analysis is general and maybe applied to other phoretic fields as well. The chemical concentration field $c(\mathbf{r},t)$ around the particles located at $\mathbf{r}_{i}$ is obtained by solving:
@@ II.2 Dynamics of the chemical field
Here, $D_{c}$ is the diffusion coefficient and $c_{0}$ is emission constant of the chemical. In the above, we have assumed that the chemical interactions between the particles is instantaneous [24, 34, 37, 38, 39, 40, 41, 42]. In addition, we assume that each particle is a point source of the chemical field. Finally, the chemical flux $\mathbf{{J}}_{i}$ on the $i$th colloid, in the point particle limit, can be written from the solution of the Eq.(3) as:
$\mathbf{J}_{i}=\frac{c_{0}}{4\pi D_{c}}\sum^{N}_{\begin{subarray}{c}j=1\\ i\neq j\end{subarray}}\frac{\mathbf{r}_{ij}}{r_{ij}^{3}}.$ (4)
Here, $r_{ij}=|\mathbf{r}_{ij}|$, with $\mathbf{r}_{ij}=\mathbf{r}_{i}-\mathbf{r}_{j}$.
Figure 2: CLF for $\Lambda_{t}=0$. (a) Phase diagram in the $(\phi,\Lambda_{r})$ plane, where $\phi$ is area fraction of particles. Here, $\Lambda_{r}^{-}$ is annotated as a dotted red line, while the density dependent $\Lambda_{r}^{+}$ in dashed black line. Symbols here and throughout the paper denoting a flock (yellow triangle), bands (green circles), and disordered (blue squares). Marked (I)-(III) scenario and transition lines (black and red) are described in panels (f-h). Snapshots of the respective phases are shown in (b)-(d) respectively (color coding for orientation of each particle is in middle panel). (e) The evolution of the polarization is displayed for the corresponding states (similarly colored). Panels (f-h) are series of scenarios depicted corresponding to parameters choices labelled in panel (a). Scenarios (I) and (III) have chemical response that are either too slow or too quick, such that a stable flock cannot be formed. Panel (g)- scenario (II) - shows the regime in which the chemical response is in an optimal range such the pair collision mechanism that would produce a flock. See III.1 for details of the three distinct cases and mechanism of flocking.
It is worthwhile to note that in our model, the chemical field diffuses in an infinite three-dimensional half-space, and thus, the chemical field decays as $1/r$. Assuming a no-flux boundary condition at the bounding surface, the $1/r$ dependence of the chemical field remains valid. On the other hand, the particles are confined to move in two-dimensions, such that they move in the plane of the bounding surface. Consequently, the dimensions of chemical concentration is: $c]=[\text{L}^{-3}]$, and chemical flux are $\mathbf{J}]=[\text{L}^{-4}]$. So, the dimensions of parameters are $\chi_{r}]=[\text{L}^{4}\text{T}^{-1}]$ and $\chi_{t}]=[\text{L}^{5}\text{T}^{-1}]$.
II.3 Important dimensionless parameters
We note that $\tau={b}/{v_{0}}$ is the propulsion time scale. In addition, we define $\tau_{r}=b^{4}/\chi_{r}$, which sets the time scale for deterministic rotation of the particle due to phoretic interactions. The phenomenology of the CLF can also be understood by studying the dynamics as per the dimensionless numbers of the problem that arise out of the equations of motion (1). They are:
$\Lambda_{r}$ $=\frac{\tau}{\tau_{r}}=\frac{\chi_{r}}{b^{3}v_{0}},\qquad\quad\Lambda_{t}=\frac{\chi_{t}}{b^{4}v_{0}}.$ (5)
Here, $\Lambda_{r}$ is the ratio of propulsion time scale $\tau$ and time scale $\tau_{r}$ which controls rotations of the particle orientations from chemical interactions between particles. On the other hand, $\Lambda_{t}$ is the ratio of translation due to chemical interactions between particles and one-body propulsion. The key parameters of this model are the dimensionless numbers $\Lambda_{r}$, $\Lambda_{t}$ and $\phi$. Here, $\phi$ is the the area fraction. $\phi$ is related to the number density $\rho$, as $\phi={\left(N\pi b^{2}\right)}/{L^{2}}=\rho\pi b^{2}$. The simulation results are henceforth presented as a function of these three parameters.
We note that in principle, there are other relevant dimensionless numbers; e.g. $\text{Pe}_{1}=\frac{v_{0}b}{D_{t}}$ and $\text{Pe}_{2}=\frac{v_{0}}{bD_{r}}$ which would alter the flocking transition. Our results are presented for effectively noiseless systems, thus the aforementioned $\text{Pe}_{1,2}\rightarrow\infty$. Indeed, the novelty of the mechanism presented here, as we will show below, is its purely deterministic destabilization of the polar phase, independent of any noise contributions.
II.4 Simulation details
The position and orientations of the particles are updated using a forward Euler-Maruyama method from the dynamical equations given in Eq. 1. Initially particles are distributed randomly over the two-dimensional space. The initial orientations are also randomly distributed over the range of angles $\pi,+\pi$] (angles are computed with respect to the positive $x$-axis). Periodic boundary conditions are applied on both the $x$-axis and $y$-axis. To compute chemical interactions using the expression of $\mathbf{J}_{i}$, we need to sum over periodic images. We have used a regular summation convention (using minimum image convention) for all results reported here, and checked these with the full Ewald summation convention for selected parameter values. In this paper, the fluctuations (noise terms) are set to be negligibly small to emphasize that flocks can be formed and destroyed by deterministic causes. Indeed, the polar order is expected to be destroyed at larger values of fluctuations [5]. But we do not explore the role of the noise in the positional and rotational sectors in this paper. Instead, we choose their strengths to be sub-dominant compared to deterministic effects due to chemical interactions and self-propulsion. We have taken particle radius $b=1$, total number of particles varies from the range $N=[239,20000]$. The steady-state time averages of the observables are taken over $2\times 10^{5}$ time steps. The time-stepping of the Euler-Maruyama integrator $dt$ is taken as $0.01$. A summary of all of the values of parameters used for the simulation are given in Table (1) in the Appendices.
III CLF via chemo-repulsive torques